Two angles are supplementary. One angle measures 12 degrees more than the other. What is the measure of the angles.

Found 2 solutions by rfer, drj:
Answer by rfer(10417)

x+x+12=180
2x+12=180
2x=168
x=84
x+12=96
96+84=180

Answer by drj(1380)

Two angles are supplementary. One angle measures 12 degrees more than the other. What is the measure of the angles.

Step 1. Supplementary angles means that two angles add up to 180 degrees.

Step 2. Let x be one angle and let x+12 be its supplementary angle and one angle measures 12 degrees more than the other

Step 3. Then x+x+12=180 or 2x+12=180. Subtracting 12 from both sides of the equation yields 2x=168. Dividing by 2 to both sides yields x=84 degrees. Then x+12=96. Also, note that 84+96 add up to 180 degrees.

Step 4. ANSWER: The supplementary angles are 84 and 96 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit
http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit
http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J